six years ago i was three times as old as my brother and now i am twice as old; find our present ages?
Answer by stanbon(75887)

six years ago i was three times as old as my brother and now i am twice as old; find our present ages?
--------------
Equations:
(i-6) = 3(b-6)
i = 2b
-------------------
Substitute and solve for "b":
(2b-6) = 3b-18
b = 12 (brother's age now)
---
Since i = 2b, i = 2*12 = 24 (my age now)
=============================================
Cheers,
Stan H.